Can someone please help me with this equation?
log(x) + log(x - 99,990) = 6
x = ______
Jerry
jerry@san.rr.com

Answer by stanbon(75887)

Can someone please help me with this equation?
log(x) + log(x - 99,990) = 6
---
log[x(x-99,990)] = 6
---
x^2-99,900x = 10*6
---
x^2 -99,900x - 1,000,000 = 0
---------
Quadratic Formula:
x = [99900 +- sqrt(99900^2 - 4000000)]/2
------
x = [99900 +- 99880]/2
----
x = 10 or x =99890
======================
Cheers,
Stan H.